OpenStudy (bossimbacon):

What is the slope of the line passing through the points (1, -5) and (4, 1)? A. 2 B. -4/5 C. 5/4 D. -2 @Destinymasha

OpenStudy (bossimbacon):

@campbell_st @ganeshie8

OpenStudy (anonymous):

\[\frac{ y2 - y1}{ x2 - x1}\]

OpenStudy (campbell_st):

just use the formula above....

OpenStudy (bossimbacon):

@shrutipande9

OpenStudy (anonymous):

If you have pick one from N answers , solution may be: 1. determine if line have positive slope (if x rise then y rise). In your case it's true. So possible good answers are 2 and 5/4. Then easy check that slope 2 is right or 5/4 is wrong. 5/4 is wrong because is 5/4 = -y1/x2 and this isn't formula for slope